THE LENGTH OF A RECTANGLE IS 6 FEET MORE THEN THE WIDTH.THE PERIMETER OF THE RECTANGLE IS 120 FEET. FIND THE LENGTH
Answer by edjones(8007)

L=w+6
P=120
2(L+w)=120
2(w+6+w)=120
2(2w+6)=120
4w+12=120
4w=108
w=27
L=33
.
Ed